# Scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes and rough fractional diffusions

### v6

@@ 4.4 Proof of Lemma 4
4.5 Proof of Lemma 6
@@
Applying Doob’s $L^{2}$-inequality to the martingale $M^{T}$, we get:
@@
$=\sum_{s=1}^{\lfloor T\rfloor}\mathbb{E}[\lambda_{s}^{T}]$ $\text{(Since }X_{s}^{T}|\mathcal{F}_{s-1}\sim\text{Poisson}(\lambda_{s}^{T})\text{)}$
@@
We now establish the asymptotic behavior of $\mathbb{E}[N_{\lfloor T\rfloor}^{T}]$. The expectation $\mathbb{E}[X_{s}^{T}]$ satisfies the discrete renewal equation $\mathbb{E}[X_{s}^{T}]=\mu^{T}+\sum_{k=1}^{s-1}\eta_{k}^{T}\mathbb{E}[X_{s-k}^{T}]$. Since $\|\eta^{T}\|_{1}=a_{T}<1$, as $s\to\infty$, $\mathbb{E}[X_{s}^{T}]$ converges to the stationary mean $\mu^{T}/(1-a_{T})$. By Cesàro’s theorem on the mean of convergent sequences, we have:
$\frac{1}{n}\mathbb{E}[N_{n}^{T}]=\frac{1}{n}\sum_{s=1}^{n}\mathbb{E}[X_{s}^{T}]\to\frac{\mu^{T}}{1-a_{T}}\quad\text{as }n\to\infty.$
Thus, for $n=\lfloor T\rfloor$, we have the asymptotic equivalence $\mathbb{E}[N_{\lfloor T\rfloor}^{T}]\sim\frac{\mu^{T}T}{1-a_{T}}$. Substituting this back, we get:
$\mathbb{E}\left[\sup_{t\in[0,1]}\left(Y_{t}^{T}-\Lambda_{t}^{T}\right)^{2}\right]$ $\leq C\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}\left(\frac{\mu^{T}T}{1-a_{T}}\right)$
$\leq C^{\prime}T^{-2\alpha}$
for some constants $C$ and $C^{\prime}$. So this term converges to 0 as $T\to\infty$. Convergence in $L^{2}$ implies convergence in probability, which completes the proof.
4.6 Proof of Lemma 7
@@
Taking the $z$-transform on both sides of equation (11) and using the convolution property Oppenheim (1999), we obtain:
@@ 4.7 Proof of Lemma 8
4.8 Proof of Lemma 9
The proof relies on the application of the discrete renewal equation from Lemma 7.
First, let’s prove Equation (19). By taking the expectation and using the law of total expectation:
@@ 4.8 Proof of Lemma 9
This is a discrete renewal equation for $x_{n}=\mathbb{E}[N_{n}^{T}]$ with driving term $y_{n}=\sum_{i=1}^{n}\hat{\mu}_{T}(i)$. Applying the solution from Lemma 7, we have:
@@
Next, for Equation (20), we start by expressing $N_{n}^{T}$ in terms of its martingale and intensity components:
@@
Applying the solution from Lemma 7 with driving term $y_{n}=M_{n}^{T}$ gives the result:
@@
5 Proof of the Main Results
5.1 Proof of Proposition 1
@@
1. Tightness of $(Z^{T},Y^{T})$. From Lemma 4, we have already established that the sequences $(Y^{T})$ and $(\Lambda^{T})$ are $\mathbb{C}$-tight. The process $Z^{T}$ is defined as
@@
Since a linear combination of $\mathbb{C}$-tight sequences is also $\mathbb{C}$-tight, $(Z^{T})$ is $\mathbb{C}$-tight. By Theorem VI-4.13 in Jacod and Shiryaev (2013), this implies that the sequence $(Z^{T})$ is tight in the Skorokhod space. As $(Y^{T})$ is also tight (from $\mathbb{C}$-tightness), the sequence of pairs $(Z^{T},Y^{T})$ is jointly tight.
2. Convergence of the Quadratic Variation. Let $(Z,Y)$ be any limit point of $(Z^{T},Y^{T})$. To prove that $Y=[Z,Z]$, we follow the standard procedure of showing that both $Z^{T},Z^{T}]_{t}$ and $Y^{T}_{t}$ converge in probability to the same process. We do this by showing that
$Z^{T},Z^{T}]_{t}-\langle Z^{T},Z^{T}\rangle_{t}\to 0$
@@
$\langle Z^{T},Z^{T}\rangle_{t}-Y^{T}_{t}\to 0.$
@@
(a) Convergence of $Z^{T},Z^{T}]_{t}$ to $\langle Z^{T},Z^{T}\rangle_{t}$. The predictable quadratic variation of $Z^{T}$ is
$\langle Z^{T},Z^{T}\rangle_{t}=$ $\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}[(\Delta Z_{s/T}^{T})^{2}|\mathcal{F}_{s-1}]$
$=$ $(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}[(X_{s}^{T}-\lambda_{s}^{T})^{2}|\mathcal{F}_{s-1}]$
$=$ $(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}.$
The sharp bracket (quadratic variation) is
@@
To show this convergence, we analyze the expectation of the sum. From the moment calculations in the proof of Lemma 4, we know there exist uniform constants $\bar{C}_{1},\bar{C}_{2}$ such that for all $s$ and $T$:
@@
For the pre-factor, by Assumption 2, there exists a constant $\bar{C}_{6}$ such that for all sufficiently large $T$:
@@
(b) Convergence of $\langle Z^{T},Z^{T}\rangle_{t}$ and $Y^{T}_{t}$ to the same limit. Our goal is to show that $\langle Z^{T},Z^{T}\rangle_{t}$ and $Y^{T}_{t}$ converge in probability to the same limit.
First, we establish a crucial identity. From the definitions of $C_{T}$ and $D_{T}$, we have the scaling factor for the martingale increments as $(C_{T}D_{T})$. Its square is:
$(C_{T}D_{T})^{2}=\left(\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}\right)\cdot\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}=D_{T}.$
Using this, the predictable quadratic variation is:
$\langle Z^{T},Z^{T}\rangle_{t}=(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}=D_{T}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}.$
By the definition of $\Lambda_{t}^{T}$, this is precisely:
$\langle Z^{T},Z^{T}\rangle_{t}=\Lambda_{t}^{T}.$
Therefore, showing that $\langle Z^{T},Z^{T}\rangle_{t}$ and $Y^{T}_{t}$ converge to the same limit is equivalent to showing that $\Lambda_{t}^{T}$ and $Y^{T}_{t}$ converge to the same limit.
From Lemma 6, we have already proven that
$\sup_{t\in[0,1]}|Y_{t}^{T}-\Lambda_{t}^{T}|\to 0\quad\text{in probability}.$
This directly implies that the two processes converge to the same limit. Since the sequence $(Y^{T})$ is tight and converges to a limit process $Y$, it follows that $\Lambda_{t}^{T}\to Y_{t}$ in probability as well. Given our identity $\langle Z^{T},Z^{T}\rangle_{t}=\Lambda_{t}^{T}$, we can conclude that
$\langle Z^{T},Z^{T}\rangle_{t}\to Y_{t}\quad\text{in probability}.$
Combining (a) and (b), we have $Z^{T},Z^{T}]_{t}\to Y_{t}$ in probability.
@@
Furthermore, we have established that $(Z^{T})$ is $\mathbb{C}$-tight. A sequence of martingales that is $\mathbb{C}$-tight and whose quadratic variation converges to a continuous process must converge to a continuous martingale (see, e.g., Proposition VI-3.26 in Jacod and Shiryaev (2013)). Since $Z^{T},Z^{T}]_{t}\to Y_{t}$ and $Y$ (as the limit of a $\mathbb{C}$-tight sequence $Y^{T}$) is continuous, the limit process $Z$ must be a continuous martingale.
5.2 Proof of Theorem 1
In fact, we only need to verify that the limit process satisfies equation (12). The proof of the Hölder property for $Y$ in Theorem 1 is exactly the same as Jaisson and Rosenbaum (2016). Like the continuous case in Jaisson and Rosenbaum (2016). We consider a converging subsequence $(Z^{T_{n}},Y^{T_{n}})$ and write $(Z,Y)$ its limit. Abusing notation slightly, we write $(Z^{T},Y^{T})$ instead of $(Z^{T_{n}},Y^{T_{n}})$. Using Skorohod’s representation theorem, there exists a probability space on which one can define copies in law of the $(Z^{T},Y^{T})$ converging almost surely under the skorohod topology to a random variable with the same law as $(Z,Y)$. We now work with this sequence of variables converging almost surely and their limit. The processes $Z$ and $Y$ being continuous, we obtain
@@ 5.2 Proof of Theorem 1
Let us now rewrite $\sum_{s=1}^{n}\lambda_{s}^{T}$. We can use Lemma 7 to obtain the expression of $\sum_{s=1}^{n}\lambda_{s}^{T}$, for all $n\in\mathbb{N}$, using Abel’s lemma, we obtain
@@
Then we try to use Lemma 7, we first calculate
$\sum_{s=1}^{n-1}A_{n-s}^{T}\sum_{u=1}^{s}\eta_{s-u}^{T}M_{u}^{T}=$ $\sum_{u=1}^{n-1}M_{u}^{T}\sum_{s=1}^{n-u-1}A_{n-s-u}^{T}\eta_{s}^{T}=\sum_{u=1}^{n-1}M_{u}^{T}(A_{s-u}^{T}-\eta_{n-u}^{T}),$
where we use the fact that $A^{T}*\eta^{T}=A^{T}-\eta^{T}$. By Lemma 7, we obtain
@@
and $\rho^{T}(\cdot)$ is defined in (10). Since $u_{T}$ converges to $1$, we get that $T_{1}\rightarrow 0$. By integrating by parts,

### v7

@@ Appendix A Proof of Lemmas
C-tightness of $(\Lambda^{T})$.
It remains to show that $(\Lambda^{T})$ is $C$-tight. Recall that by Lemma 2.9,
$\sup_{t\in[0,1]}\left|Y_{t}^{T}-\Lambda_{t}^{T}\right|\xrightarrow[T\to\infty]{\mathbb{P}}0.$ (A.5)
Since $(Y^{T})$ is $C$-tight, for any $\varepsilon>0$ there exists a compact set $K\subset C([0,1])$ such that $\inf_{T}\mathbb{P}(Y^{T}\in K)\geq 1-\varepsilon$. For $\delta>0$, let
$K^{\delta}:=\left\{x\in D([0,1]):\ \inf_{y\in K}\|x-y\|_{\infty}\leq\delta\right\},$
which is relatively compact in $D([0,1])$. Then, for all $T$,
$\mathbb{P}(\Lambda^{T}\in K^{\delta})\geq\mathbb{P}\left(Y^{T}\in K,\ \|Y^{T}-\Lambda^{T}\|_{\infty}\leq\delta\right)\geq 1-\varepsilon-\mathbb{P}\left(\|Y^{T}-\Lambda^{T}\|_{\infty}>\delta\right).$
Letting $T\to\infty$ and using (A.5) yields $\liminf_{T\to\infty}\mathbb{P}(\Lambda^{T}\in K^{\delta})\geq 1-\varepsilon$. This proves tightness of $(\Lambda^{T})$. Moreover, any limit point of $(\Lambda^{T})$ must coincide with the (continuous) limit of $(Y^{T})$ because of (A.5), hence $(\Lambda^{T})$ is in fact $C$-tight.
A.5 Proof of Lemma 2.7
Proof.
We first note that for each $T$ and $m\geq 1$,
$(\eta^{T*m})_{k}=(a_{T}\eta)^{*m}_{k}=a_{T}^{m}(\eta^{*m})_{k},$
and therefore
$A_{k}^{T}=\sum_{m\geq 1}(\eta^{T*m})_{k}=\sum_{m\geq 1}a_{T}^{m}(\eta^{*m})_{k}\leq\sum_{m\geq 1}(\eta^{*m})_{k}=A_{k},$
since $a_{T}\in(0,1)$ and all terms are nonnegative. Hence it suffices to prove that $A_{k}\leq Ck^{\alpha-1}$.
To this end, define the Laplace (power-series) transforms
$\hat{\eta}(s):=\sum_{n\geq 1}\eta_{n}e^{-sn},\qquad s>0,\qquad\text{and}\qquad\hat{A}(s):=\sum_{k\geq 1}A_{k}e^{-sk},\qquad s>0.$
Because $\eta$ is a probability mass function, $\hat{\eta}(s)\in(0,1)$ for every $s>0$. Using the convolution property and Fubini’s theorem (all terms are nonnegative), we obtain
$\hat{A}(s)$ $=\sum_{k\geq 1}\sum_{m\geq 1}(\eta^{*m})_{k}e^{-sk}=\sum_{m\geq 1}\sum_{k\geq 1}(\eta^{*m})_{k}e^{-sk}=\sum_{m\geq 1}\left(\sum_{n\geq 1}\eta_{n}e^{-sn}\right)^{m}$
$=\sum_{m\geq 1}\hat{\eta}(s)^{m}=\frac{\hat{\eta}(s)}{1-\hat{\eta}(s)}.$
Next we analyze $1-\hat{\eta}(s)$ as $s\downarrow 0$. By the same Abel-summation argument as in the proof of Lemma 2.4,
$\hat{\eta}(s)=1-(1-e^{-s})\sum_{n=0}^{\infty}e^{-sn}\sum_{k=n+1}^{\infty}\eta_{k}.$
Under Assumption 2.1(i), the tail satisfies
$\sum_{k=n+1}^{\infty}\eta_{k}\sim\frac{K}{\alpha}\,n^{-\alpha}\qquad(n\to\infty).$
Applying Lemma 2.5 with
$\beta_{n}:=\sum_{k=n+1}^{\infty}\eta_{k},\quad\rho:=-\alpha,\quad c:=\frac{K}{\alpha}\Gamma(1-\alpha),$
(which is legitimate since $\alpha\in(\frac{1}{2},1)$ implies $\rho=-\alpha>-1$) yields
$(1-e^{-s})\sum_{n=0}^{\infty}e^{-sn}\sum_{k=n+1}^{\infty}\eta_{k}\sim\frac{K}{\alpha}\Gamma(1-\alpha)\,s^{\alpha}=\delta\,s^{\alpha}\qquad(s\downarrow 0),$
where $\delta:=K\Gamma(1-\alpha)/\alpha$. Consequently,
$1-\hat{\eta}(s)\sim\delta\,s^{\alpha}\qquad(s\downarrow 0).$
Combining this with $\hat{\eta}(s)\to 1$ as $s\downarrow 0$, we deduce
$\hat{A}(s)=\frac{\hat{\eta}(s)}{1-\hat{\eta}(s)}\sim\frac{1}{\delta}\,s^{-\alpha}\qquad(s\downarrow 0).$
Now define the partial sums
$U_{n}:=\sum_{k=1}^{n}A_{k},\qquad n\geq 1.$
A summation-by-parts identity gives
$(1-e^{-s})\sum_{n=0}^{\infty}e^{-sn}U_{n}=\sum_{n=0}^{\infty}(e^{-sn}-e^{-s(n+1)})U_{n}=\sum_{k\geq 1}A_{k}e^{-sk}=\hat{A}(s),$
and hence
$(1-e^{-s})\sum_{n=0}^{\infty}e^{-sn}U_{n}\sim\frac{1}{\delta}\,s^{-\alpha}\qquad(s\downarrow 0).$
Since $(U_{n})$ is nondecreasing, we may apply the discrete Karamata Tauberian theorem for power series (e.g. (Bingham et al., 1989, Corollary 1.7.3)) to obtain
$U_{n}\sim\frac{1}{\delta\,\Gamma(1+\alpha)}\,n^{\alpha}\qquad(n\to\infty).$
In particular, there exist $C_{0}>0$ and $n_{0}$ such that $U_{n}\leq C_{0}n^{\alpha}$ for all $n\geq n_{0}$. Therefore, for $n\geq n_{0}$,
$A_{n}=U_{n}-U_{n-1}\leq C_{0}\left(n^{\alpha}-(n-1)^{\alpha}\right)\leq C_{0}\,\alpha\,n^{\alpha-1},$
where we used the mean-value theorem (equivalently, the concavity of $x\mapsto x^{\alpha}$ on $(0,\infty)$). Enlarging the constant to also cover $1\leq n<n_{0}$ yields
$A_{n}\leq C\,n^{\alpha-1}\qquad\text{for all }n\geq 1.$
Finally, since $A_{k}^{T}\leq A_{k}$ for all $k$, we conclude that
$A_{k}^{T}\leq A_{k}\leq C\,k^{\alpha-1},\qquad\forall\,T,\ \forall\,k\geq 1,$
which completes the proof. ∎
A.6 Proof of Lemma 2.9
@@
Applying Doob’s $L^{2}$-inequality to the martingale $(Y_{t}^{T}-\Lambda_{t}^{T})_{t\in[0,1]}$ (equivalently to $M^{T}$), we get:
@@
$=\sum_{s=1}^{\lfloor T\rfloor}\mathbb{E}[\lambda_{s}^{T}]$ $\text{(Since }X_{s}^{T}|\mathcal{F}_{s-1}\sim\text{Pois}(\lambda_{s}^{T})\text{)}$
@@
We now bound $\mathbb{E}[N_{\lfloor T\rfloor}^{T}]$. Using the tower property and $\mathbb{E}[X_{s}^{T}\mid\mathcal{F}_{s-1}]=\lambda_{s}^{T}$, we have
$\mathbb{E}[N_{\lfloor T\rfloor}^{T}]=\sum_{s=1}^{\lfloor T\rfloor}\mathbb{E}[X_{s}^{T}]=\sum_{s=1}^{\lfloor T\rfloor}\mathbb{E}[\lambda_{s}^{T}]\leq\lfloor T\rfloor\sup_{1\leq s\leq\lfloor T\rfloor}\mathbb{E}[\lambda_{s}^{T}].$
By the uniform first-moment bound (A.3), there exists $\bar{C}_{1}>0$ such that $\sup_{1\leq s\leq\lfloor T\rfloor}\mathbb{E}[\lambda_{s}^{T}]\leq\bar{C}_{1}T^{2\alpha-1}$ for all $T\geq 1$. Therefore,
$\mathbb{E}[N_{\lfloor T\rfloor}^{T}]\leq\bar{C}_{1}\,T\cdot T^{2\alpha-1}=\bar{C}_{1}T^{2\alpha}.$
Substituting this bound back, we get:
$\mathbb{E}\left[\sup_{t\in[0,1]}\left(Y_{t}^{T}-\Lambda_{t}^{T}\right)^{2}\right]$ $\leq 4\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}\mathbb{E}\left[(M_{\lfloor T\rfloor}^{T})^{2}\right]$
$=4\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}\mathbb{E}\left[N_{\lfloor T\rfloor}^{T}\right]$
$\leq 4\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}\cdot\bar{C}_{1}T^{2\alpha}\;\leq\;C^{\prime}(1-a_{T})^{2},$
for some constant $C^{\prime}>0$. Since $a_{T}\to 1$ by Assumption 2.3, we have $(1-a_{T})^{2}\to 0$, hence the right-hand side converges to 0.
A.7 Proof of Lemma 2.11
@@
Taking the $z$-transform on both sides of equation (2.9) and using the convolution property Oppenheim (1999), we obtain:
@@
A.9 Proof of Lemma 2.17
The proof relies on the application of the discrete renewal equation from Lemma 2.11.
First, let’s prove Equation (2.15). By taking the expectation and using the law of total expectation:
@@
This is a discrete renewal equation for $x_{n}=\mathbb{E}[N_{n}^{T}]$ with driving term $y_{n}=\sum_{i=1}^{n}\hat{\mu}_{T}(i)$. Applying the solution from Lemma 2.11, we have:
@@
Next, for Equation (2.16), we start by expressing $N_{n}^{T}$ in terms of its martingale and intensity components:
@@
Applying the solution from Lemma 2.11 with driving term $y_{n}=M_{n}^{T}$ gives the result:
@@
Appendix B Proof of the Main Results
B.1 Proof of Proposition 2.10
@@ Appendix B Proof of the Main Results
1. Tightness of $(Z^{T},Y^{T})$. From Lemma 2.6, we have already established that the sequences $(Y^{T})$ and $(\Lambda^{T})$ are $C$-tight. The process $Z^{T}$ is defined as
@@
Since $(Y^{T})$ and $(\Lambda^{T})$ are $C$-tight in $\mathbb{D}([0,1])$, the pair $(Y^{T},\Lambda^{T})$ is $C$-tight in the product space $\mathbb{D}([0,1])^{2}$. Moreover, $\mathbb{D}([0,1])$ endowed with the Skorokhod $J_{1}$ topology is a topological vector space, hence the subtraction map
$\Phi:\mathbb{D}([0,1])^{2}\to\mathbb{D}([0,1]),\qquad\Phi(x,y)=x-y,$
is continuous (see, e.g., (Jacod and Shiryaev, 2013, Chapter VI, §1)). Therefore, by the continuous mapping theorem, $(Y^{T}-\Lambda^{T})$ is $C$-tight in $\mathbb{D}([0,1])$.
Finally, deterministic scalar multiplication is continuous on $\mathbb{D}([0,1])$, so the rescaled process
$Z_{t}^{T}=\sqrt{\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}}\,(Y_{t}^{T}-\Lambda_{t}^{T})$
is also $C$-tight.
We now prove tightness of $(Z^{T})$. Recall that
$Z_{t}^{T}=\sqrt{\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}}\,(Y_{t}^{T}-\Lambda_{t}^{T})=\sqrt{\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}}\;M_{\lfloor tT\rfloor}^{T},$
where $M_{n}^{T}=\sum_{s=1}^{n}(X_{s}^{T}-\lambda_{s}^{T})$ is the compensated martingale. Set
$D_{T}:=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}.$
Then the jump at time $s/T$ is
$\Delta Z_{s/T}^{T}=\sqrt{D_{T}}\,(X_{s}^{T}-\lambda_{s}^{T}).$
Using that $X_{s}^{T}\mid\mathcal{F}_{s-1}\sim\mathrm{Pois}(\lambda_{s}^{T})$, the conditional fourth centered moment satisfies
$\mathbb{E}\!\left[(X_{s}^{T}-\lambda_{s}^{T})^{4}\mid\mathcal{F}_{s-1}\right]=\lambda_{s}^{T}+3(\lambda_{s}^{T})^{2},$
hence, by the moment bounds on $\lambda_{s}^{T}$ proved in Lemma 2.6,
$\sup_{1\leq s\leq\lfloor T\rfloor}\mathbb{E}\!\left[(X_{s}^{T}-\lambda_{s}^{T})^{4}\right]\leq C\,T^{4\alpha-2}\qquad(T\geq 1).$
Therefore, by a union bound and Markov’s inequality,
$\mathbb{P}\!\left(\sup_{t\in[0,1]}|\Delta Z_{t}^{T}|>\varepsilon\right)\leq$ $\sum_{s=1}^{\lfloor T\rfloor}\frac{\mathbb{E}[(\Delta Z_{s/T}^{T})^{4}]}{\varepsilon^{4}}$
$\leq$ $\frac{T\,D_{T}^{2}}{\varepsilon^{4}}\,\sup_{s\leq\lfloor T\rfloor}\mathbb{E}[(X_{s}^{T}-\lambda_{s}^{T})^{4}]$
$\leq$ $\frac{C}{\varepsilon^{4}}\,T\cdot T^{-4\alpha}\cdot T^{4\alpha-2}=\frac{C}{\varepsilon^{4}}\,T^{-1}\to 0.$
In particular, $\sup_{t\in[0,1]}|\Delta Z_{t}^{T}|\xrightarrow{\mathbb{P}}0$.
Next, the predictable quadratic variation of $Z^{T}$ is
$\langle Z^{T}\rangle_{t}=\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}\!\left[(\Delta Z_{s/T}^{T})^{2}\mid\mathcal{F}_{s-1}\right]=D_{T}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}=\Lambda_{t}^{T}.$
Since $(\Lambda^{T})$ is $C$-tight by Lemma 2.6, the family $(\langle Z^{T}\rangle)$ is $C$-tight in $\mathbb{D}([0,1])$ and any limit point is a.s. continuous.
Moreover, $\sup_{t\in[0,1]}|\Delta Z_{t}^{T}|\xrightarrow{\mathbb{P}}0$, with $\Delta Z_{t}^{T}:=Z_{t}^{T}-Z_{t-}^{T}$, i.e. the ARJ(1)(1) condition of Rebolledo (1980). Therefore, by (Rebolledo, 1980, Proposition 12), $(Z^{T})$ is $C$-tight in $\mathbb{D}([0,1])$.
Finally, since both $(Z^{T})$ and $(Y^{T})$ are tight in $\mathbb{D}([0,1])$, the pair $(Z^{T},Y^{T})$ is tight in $\mathbb{D}([0,1])^{2}$.
2. Convergence of the Quadratic Variation. Let $(Z,Y)$ be any limit point of $(Z^{T},Y^{T})$. To prove that $Y=[Z,Z]$, we follow the standard procedure of showing that both $Z^{T},Z^{T}]$ and $Y^{T}$ have the same limit. More precisely, we prove the following two convergences in probability, uniformly on $[0,1]$:
$\sup_{t\in[0,1]}\left|[Z^{T},Z^{T}]_{t}-\langle Z^{T},Z^{T}\rangle_{t}\right|\xrightarrow[T\to\infty]{\mathbb{P}}0,$}
@@
$\sup_{t\in[0,1]}\left|\langle Z^{T},Z^{T}\rangle_{t}-Y^{T}_{t}\right|\xrightarrow[T\to\infty]{\mathbb{P}}0.$}
In particular, for each fixed $t\in[0,1]$, these convergences hold in probability at time $t$.
@@
(a) Asymptotic equivalence of $Z^{T},Z^{T}]$ and $\langle Z^{T},Z^{T}\rangle$. Note that both $Z^{T},Z^{T}]$ and $\langle Z^{T},Z^{T}\rangle$ depend on $T$; what we prove is that their difference is asymptotically negligible:
$\sup_{t\in[0,1]}\left|[Z^{T},Z^{T}]_{t}-\langle Z^{T},Z^{T}\rangle_{t}\right|\xrightarrow[T\to\infty]{\mathbb{P}}0.$
The predictable quadratic variation of $Z^{T}$ is given by the sum of conditional second moments of its increments.
$\langle Z^{T},Z^{T}\rangle_{t}=$ $\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}\!\left[\left.(\Delta Z_{s/T}^{T})^{2}\,\right|\,\mathcal{F}_{s-1}\right]$
$=$ $(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}\!\left[\left.(X_{s}^{T}-\lambda_{s}^{T})^{2}\,\right|\,\mathcal{F}_{s-1}\right]$
$=$ $(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\mathrm{Var}\!\left(\left.X_{s}^{T}\,\right|\,\mathcal{F}_{s-1}\right)$
$=$ $(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T},$
where we used that, by construction, $X_{s}^{T}\mid\mathcal{F}_{s-1}\sim\mathrm{Pois}(\lambda_{s}^{T})$, hence $\mathrm{Var}(X_{s}^{T}\mid\mathcal{F}_{s-1})=\lambda_{s}^{T}$. The sharp bracket (quadratic variation) is
@@
To show this convergence, we analyze the expectation of the sum. From the moment calculations in the proof of Lemma 2.6, we know there exist uniform constants $\bar{C}_{1},\bar{C}_{2}$ such that for all $s$ and $T$:
@@
For the pre-factor, by Assumption 2.3, there exists a constant $\bar{C}_{6}$ such that for all sufficiently large $T$:
@@
(b) Asymptotic equivalence of $\langle Z^{T},Z^{T}\rangle$ and $Y^{T}$ (same limit points). Our goal is to show that the difference between $\langle Z^{T},Z^{T}\rangle$ and $Y^{T}$ vanishes in probability, uniformly on $[0,1]$:
$\sup_{t\in[0,1]}\left|\langle Z^{T},Z^{T}\rangle_{t}-Y_{t}^{T}\right|\xrightarrow[T\to\infty]{\mathbb{P}}0.$
Equivalently, $\langle Z^{T},Z^{T}\rangle$ and $Y^{T}$ have the same limit points in $\mathbb{D}([0,1])$.
First, we establish a crucial identity. From the definitions of $C_{T}$ and $D_{T}$, we have
$(C_{T}D_{T})^{2}=\left(\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}\right)\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}=D_{T}.$
Hence the predictable quadratic variation is
$\langle Z^{T},Z^{T}\rangle_{t}=\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}\!\left[(\Delta Z_{s/T}^{T})^{2}\mid\mathcal{F}_{s-1}\right]=(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}=D_{T}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}=\Lambda_{t}^{T}.$
Therefore,
$\sup_{t\in[0,1]}\left|\langle Z^{T},Z^{T}\rangle_{t}-Y_{t}^{T}\right|=\sup_{t\in[0,1]}\left|\Lambda_{t}^{T}-Y_{t}^{T}\right|.$
By Lemma 2.9, we already have
$\sup_{t\in[0,1]}\left|\Lambda_{t}^{T}-Y_{t}^{T}\right|\xrightarrow[T\to\infty]{\mathbb{P}}0,$
and thus
$\sup_{t\in[0,1]}\left|\langle Z^{T},Z^{T}\rangle_{t}-Y_{t}^{T}\right|\xrightarrow[T\to\infty]{\mathbb{P}}0,$
which proves the claimed asymptotic equivalence (and in particular that $\langle Z^{T}\rangle$ and $Y^{T}$ share the same limit points).
@@
Furthermore, we have established that $(Z^{T})$ is $C$-tight. A sequence of martingales that is $C$-tight and whose quadratic variation converges to a continuous process must converge to a continuous martingale (see, e.g., Proposition VI-3.26 in Jacod and Shiryaev (2013)). Since $Z^{T},Z^{T}]_{t}\to Y_{t}$ and $Y$ (as the limit of a $C$-tight sequence $Y^{T}$) is continuous, the limit process $Z$ must be a continuous martingale.
B.2 Proof of Theorem 2.13
In fact, we only need to verify that the limit process satisfies equation (2.10). The proof of the Hölder property for $Y$ in Theorem 2.13 is exactly the same as in Jaisson and Rosenbaum (2016). Let $(Z^{T_{n}},Y^{T_{n}})$ be a subsequence converging in distribution to $(Z,Y)$, and for simplicity write $(Z^{T},Y^{T})$ instead of $(Z^{T_{n}},Y^{T_{n}})$. By Skorohod’s representation theorem, there exists a probability space on which one can define copies (with the same laws) of $(Z^{T},Y^{T})$ and $(Z,Y)$ such that $(Z^{T},Y^{T})\to(Z,Y)$ almost surely in the Skorohod topology. We work on this auxiliary probability space from now on. In particular, to apply the Dambis–Dubins–Schwarz theorem below, we may pass to the usual extension of this probability space so that there exists a Brownian motion $B$ such that $Z_{t}=B_{Y_{t}}$. Since the processes $Z$ and $Y$ are continuous, we obtain
@@
Let us now rewrite $\sum_{s=1}^{n}\lambda_{s}^{T}$. We can use Lemma 2.11 to obtain the expression of $\sum_{s=1}^{n}\lambda_{s}^{T}$, for all $n\in\mathbb{N}$, using Abel’s lemma, we obtain
@@
Then we try to use Lemma 2.11, we first calculate
$\sum_{s=1}^{n-1}A_{n-s}^{T}\sum_{u=1}^{s}\eta_{s-u}^{T}M_{u}^{T}=$ $\sum_{u=1}^{n-1}M_{u}^{T}\sum_{s=1}^{n-u-1}A_{n-s-u}^{T}\eta_{s}^{T}=\sum_{u=1}^{n-1}M_{u}^{T}(A_{n-u}^{T}-\eta_{n-u}^{T}),$
where we use the fact that $A^{T}*\eta^{T}=A^{T}-\eta^{T}$. By Lemma 2.11, we obtain
@@
and $\rho^{T}(\cdot)$ is defined in (2.8). As $T\to\infty$, since $u_{T}\to 1$ and $(1-a_{T})\to 0$, we have $T_{1}\to 0$. For $T_{2}$, recognized as a Riemann sum converging to a convolution integral, and by integrating by parts: